A ball on a string moves around a complete circle, once a second, on a frictionless, horizontal table. The tension in the string is measured to be 5.9 N. What would the tension be (in N) if the ball went around in 0.21 s?
Answer #1

According to the given problem,

Using the relation,

v ∝ 1/t

v1/v2 = t2/t1

v1/v2 = 0.21

Now, using the concept of centrepital force,

T = mv2/r

T ∝ v2

T1/T2 = (v1/v2)2

T2 = T1/0.212

T2 = 5.9/0.212

T2 = 133.786 = 134 N

I hope you understood the problem, If yes rate me! or else fell free to comment.